# Generalized Linear Bandits with Limited Adaptivity

### v4

@@ B.2 Confidence Sets for non-Switching Criterion I rounds
Recall that ${\mathbf{H}}_{t}$ is defined in line 15 of Algorithm 2. Further, we define
@@
Recall that $\tau$ is the round when the Switching Criterion II (Line 9) is satisfied. Now we define the following quantities:
$g_{\tau}(\theta)$ $=\sum_{s\in[\tau-1]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\theta\rangle\right)x_{s}+\lambda_{\tau}\theta$
@@
$\tilde{\theta}$ $=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})$
$\beta$ $\coloneqq\textbf{c}\sqrt{d\log(T/\delta)}$
@@
$\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{\tau}}$ $\leq 2(1+2)\left\lVert g_{\tau}(\theta^{*})-\sum_{s\in[\tau-1]}r_{s}x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}$ (by Lemma E.1)
$\leq 6\textbf{c}\sqrt{d\log(T/\delta)}$ (by Lemma B.1)
@@ B.3 Bounding the instantaneous regret
In this subsection, we will only consider rounds $t\in[T]$ which does not satisfy Switching Criterion I. Let $x_{t}\in{\mathcal{X}}_{t}$ be the played arm defined via line 14 Algorithm 2. Further, let $x^{*}_{t}\in{\mathcal{X}}_{t}$ be the best available arm in that round.
@@
$\leq\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle+\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{\tau}}\right)-\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle-\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{\tau}}\right)$ (optimistic $x_{t}$, see line 14 algo. 2)
@@
The arm set ${\mathcal{X}}^{\prime}_{t}$ obtained after eliminating arms from ${\mathcal{X}}_{t}$ (line 13 Algorithm 2), under event ${\mathcal{E}}_{w}$, satisfies: (a) $x^{*}_{t}\in{\mathcal{X}}_{t}$, (b) $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq\frac{4}{R}$
@@
Given $\delta\in(0,1)$, with probability $\geq 1-\delta$, the regret of RS-GLinCB (Algorithm 2) satisfies $\textsc{R}_{T}=O\big{(}d\sqrt{\sum_{t\in[T]}\dot{\mu}\left(\langle x_{t}^{*},\theta^{*}\rangle\right)}\log\left({{RT}/{\delta}}\right)+\ \kappa d^{2}R^{5}S^{2}\log^{2}\left({{T}/{\delta}}\right)\big{)}.$
@@
$\leq\textbf{c}R^{3}\kappa\gamma^{2}\log(T/\delta)+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}(z)\langle x^{*}_{t}-x_{t},\theta^{*}\rangle$ (some $z\in[\langle x_{t},\theta^{*}\rangle,\langle x_{t}^{*},\theta^{*}\rangle]$; lemma B.11)
@@
$\leq\textbf{c}d\log(RT/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}.$
@@
$\textsc{R}_{T}\leq\textbf{c}d\log(RT/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}+\textbf{c}R^{5}S^{2}\kappa\log(T/\delta)^{2}.$
@@
$\tau_{o}\leq 2dR^{2}\kappa\gamma^{2}\log{\left(1+\frac{\tau_{o}}{\lambda d}\right)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$
Lemma (4.1).
Algorithm 2, during its entire execution, updates its policy at most $O(R^{4}S^{2}\ \kappa d^{2}\ \log^{2}(T/\delta))$ times.
@@
Note that in Algorithm 2, policy changes happen only in the rounds when either of the Switching Criteria are triggered. The number of times Switching Criterion I is triggered is bounded by Lemma B.11. On the other hand, the number of times Switching Criterion II is triggered is equal to the number of times determinant of ${\mathbf{H}}_{t}$ doubles, which is bounded by Lemma B.15. Thus in total, the number of policy changes in Algorithm 2 is upper bounded by $2dR^{2}\kappa\gamma^{2}\log(T/\delta)+\textbf{c}d\log(T)~{}.$
@@ B.6 Some Useful Lemmas
Lemma B.12 (Elliptic Potential Lemma (Lemma 10 Abbasi-Yadkori et al. [2011])).
@@
Lemma B.13 (Lemma 12 of Abbasi-Yadkori et al. [2011]).
@@
Lemma B.14 (Lemma 10 of Abbasi-Yadkori et al. [2011]).
@@ Appendix C Useful Properties of GLMs
As a consequence, we have the following simple modification of the self-concordance results of Faury et al. [2020].
@@
Putting $z=z_{2}$ establishes 26. To show 25, we further set $z=z_{1}+u(z_{2}-z_{1})$ for $u\in[0,1]$, (note that $z\geq z_{1}$) and integrate on $u$,
@@
For 27, we have, by application of 26, $\dot{\mu}(z_{1}+v(z_{2}-z_{1}))\geq\dot{\mu}(z_{1})\exp(R|v(z_{2}-z_{1})|)$. Therefore,
@@
$\geq\dot{\mu}(z_{1})\cdot\frac{1}{2+R|z_{1}-z_{2}|}$ (Lemma 10 of Abeille et al. [2021])
@@
In Faury [2021] Section 1.4.1, the author claims that if the GLM parameter $z$ lies in a bounded set, then the GLM is self-concordant, i.e., $|\ddot{\mu}(z)|\leq a\dot{\mu}(z)$, for some appropriate constant $a$ over this bounded set. Thereafter the author notes that the techniques developed in Faury [2021] guarantees $\kappa$-free regret rates (in $\sqrt{T}$ term) for such GLMs (i.e., all GLMs with bounded parameter). However, the claim regarding self-concordance of GLMs is not true in general. There are classes of GLMs whose parameters may be restricted in a bounded set, but for them no constant $a$ exists. One such example is the exponential distribution. The link function $\mu$ for exponential distribution is given as $\mu(z)=-\frac{1}{z}$. If we allow $z$ to lie in the set $(-c,0)$ for some positive $c$, then we have $\mu(z)$ strictly increasing (satisfying our assumption on monotonicity of $\mu$, thus a valid example). However, for this GLM,
@@ Appendix D Computational Cost
Consider a log-loss minimization oracle that returns the unconstrained MLE for a given GLM class with a computational complexity of $C_{opt}\cdot n$, when the log-loss is computed over $n$ data points. Let the maximum number of arms available every round be $K$. Furher, let the computational cost of an oracle that solves the non-convex optimization 8 be $NC_{opt}$.
@@
Comparison with ECOLog Faury et al. [2022] shows that execution time for RS-GLinCB is significantly smaller. We posit that this is because RS-GLinCB solves a large convex optimization problem but less frequently, resulting into smaller overhead at the implementation level, while ECOLog solves a smaller convex optimization problem, but does so every round. On an implementation level, this translates into more function calls and computation. Further, we observe that with increasing $\kappa$, the execution time of RS-GLinCB increases, which is in accordance with Lemma 4.1 that quantifies the number of policy switches as an increasing function of $\kappa$.
While comparing with GLOC Jun et al. [2017], we observe that RS-GLinCB performs better than GLOC in both high and low $\kappa$ regimes. Since GLOC runs an online convex optimization (online Newton step) algorithm to generate its confidence sets, the time taken by GLOC is nearly constant with changing $\kappa$. On the other hand, in accordance with Lemma 4.1, the computational cost of RS-GLinCB increases with $\kappa$. However, after a few initial rounds, when neither of the switching criteria are triggered, RS-GLinCB does not need to solve any computationally intensive optimization problem, hence these rounds execute very fast. In practice, with typical data distribution, RS-GLinCB reaches this stage much before what the worst-case guarantees show, hence we see it perform better than GLOC.
@@ Appendix E Projection
We describe the projection step used in Algorithms 1 and 2. We present arguments similar to the ones made in Appendix B.3 of Faury et al. [2020]. We write
@@
$\widetilde{\theta}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}\left\lVert\sum_{s=1}^{t}(\mu\left(\langle x_{s},\theta\rangle\right)-\mu\left(\langle x_{s},\widehat{\theta}\rangle\right))x_{s}\right\rVert_{{\mathbf{H}}(\theta)^{-1}}$ (28)
@@
$\left\lVert\widetilde{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}(\theta^{*})}\leq 2(1+c)\left\lVert\sum_{s=1}^{t}(\mu\left(\langle x_{s},\theta^{*}\rangle\right)-r_{s})x_{s}\right\rVert_{{\mathbf{H}}(\theta^{*})^{-1}}$
@@
Let us define the matrix ${\mathbf{G}}=\sum_{s\in[t]}\alpha(x,\widetilde{\theta},\theta^{*})x_{s}x_{s}^{\intercal}$. Using the above fact, we obtain the relation: ${\mathbf{G}}\succeq\frac{1}{1+c}{\mathbf{H}}^{*}$ and ${\mathbf{G}}\succeq\frac{1}{1+c}{{\mathbf{H}}}(\widetilde{\theta})$. Also define the vector $g(\theta)=\sum_{s\in[t]}\mu\left(\langle\theta,x_{s}\rangle\right)x_{s}$. Now,
@@
$=\sqrt{1+c}\left\lVert\sum_{s\in[t]}\left(\alpha(x_{s},\widetilde{\theta},\theta^{*})\langle\widetilde{\theta}-\theta^{*},x_{s}\rangle\right)x_{s}\right\rVert_{{\mathbf{G}}^{-1}}$
$=\sqrt{1+c}\left\lVert\sum_{s\in[t]}\left(\mu\left(\langle x_{s},\widetilde{\theta}\rangle\right)-\mu\left(\langle x_{s},\theta^{*}\rangle\right)\right)x_{s}\right\rVert_{{\mathbf{G}}^{-1}}$ (Taylor’s theorem)
$=\sqrt{1+c}\left\lVert\left(\sum_{s\in[t]}\mu\left(\langle\widetilde{\theta},x_{s}\rangle\right)x_{s}\right)-\left(\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
Let $g(\theta)=\sum_{s=1}^{t}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}$ for any $\theta$. Therefore, we have,
@@
$=\sqrt{1+c}\left\lVert g(\widetilde{\theta})-g(\widehat{\theta})+g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}$
@@
$\leq(1+c)\left(\left\lVert g(\widetilde{\theta})-g(\widehat{\theta})\right\rVert_{{{\mathbf{H}}(\widetilde{\theta})}^{-1}}+\left\lVert g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{H}}^{*-1}}\right)$ (${\mathbf{H}}^{*-1}\succeq(\sqrt{1+c}){\mathbf{G}}^{-1}$)
$\leq 2(1+c)\left\lVert g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{H}}^{*-1}}$ (by (28))
$=2(1+c)\left\lVert g(\theta^{*})-\sum_{s\in[t]}r_{s}x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}$ ($\widehat{\theta}$ is the unconstrained MLE, $g(\widehat{\theta})=\sum_{s\in[t]}r_{s}x_{s}$.)

### v5

@@ B.2 Confidence Sets for non-Switching Criterion I rounds
Recall that ${\mathbf{H}}_{t}$ is defined in line 16 of Algorithm 2. Further, we define
@@
Recall that $\tau$ is the round when the Switching Criterion II (Line 10) is satisfied. Now we define the following quantities:
$g_{\tau}(\theta)$ $=\sum_{s\in[\tau-1]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\theta\rangle\right)x_{s}+\lambda\theta$
@@
$\tilde{\theta}$ $=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\lambda\left\lVert\theta\right\rVert_{2}^{2}$
$\beta$ $\coloneqq\textbf{c}R\sqrt{dS\log(ST/\delta)}$
@@
$\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{\tau}}$ $\leq 2\left(1+2\right)\left\lVert\sum_{s\in[\tau-1]\setminus{\mathcal{T}}_{o}}\left(r_{s}-\mu\left(\langle x_{s},\theta^{*}\rangle\right)\right)x_{s}\right\rVert_{{\mathbf{H}}_{\tau}^{*-1}}+2S\sqrt{\lambda\left(1+2\right)}$ (by Lemma E.1)
$\leq\textbf{c}R\sqrt{dS\log(ST/\delta)}$ (by Lemma B.1)
@@ B.3 Bounding the instantaneous regret
In this subsection, we will only consider rounds $t\in[T]$ which does not satisfy Switching Criterion I. Let $x_{t}\in{\mathcal{X}}_{t}$ be the played arm defined via line 15 Algorithm 2. Further, let $x^{*}_{t}\in{\mathcal{X}}_{t}$ be the best available arm in that round.
@@
$\leq\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle+\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{\tau}}\right)-\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle-\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{\tau}}\right)$ (optimistic $x_{t}$, see line 15 algo. 2)
@@
The arm set ${\mathcal{X}}^{\prime}_{t}$ obtained after eliminating arms from ${\mathcal{X}}_{t}$ (line 14 Algorithm 2), under event ${\mathcal{E}}_{w}$, satisfies: (a) $x^{*}_{t}\in{\mathcal{X}}_{t}$, (b) $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq\frac{4}{R}$
@@
Given $\delta\in(0,1)$, with probability $\geq 1-\delta$, the regret of RS-GLinCB (Algorithm 2) satisfies $\textsc{R}_{T}=O\big{(}dRS^{\nicefrac{{1}}{{2}}}\sqrt{\sum_{t\in[T]}\dot{\mu}\left(\langle x_{t}^{*},\theta^{*}\rangle\right)}\log\left({{ST}/{\delta}}\right)+\ \kappa d^{2}R^{7}S^{3}\log^{2}\left({{ST}/{\delta}}\right)\big{)}.$
@@
$\leq 2dR^{3}\kappa\gamma^{2}\log{(ST/\delta)}+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}(z)\langle x^{*}_{t}-x_{t},\theta^{*}\rangle$ (some $z\in[\langle x_{t},\theta^{*}\rangle,\langle x_{t}^{*},\theta^{*}\rangle]$; lemma B.12)
@@
$\leq\textbf{c}RS^{\nicefrac{{1}}{{2}}}d\log(ST/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}.$
@@
$\textsc{R}_{T}\leq\textbf{c}RS^{\nicefrac{{1}}{{2}}}d\log(ST/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}+\textbf{c}R^{7}S^{3}\kappa d^{2}\log(ST/\delta)^{2}.$
@@
$\tau_{o}\leq 2dR^{2}\kappa\gamma^{2}\log{\left(1+\frac{\tau_{o}}{\lambda d}\right)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T)}\leq 2dR^{2}\kappa\gamma^{2}\log{(ST/\delta)}$
Lemma B.13.
Algorithm 2, during its entire execution, updates its policy at most $O(R^{6}S^{3}\ \kappa d^{2}\ \log^{2}(T/\delta))$ times.
@@
Note that in Algorithm 2, policy changes happen only in the rounds when either of the Switching Criteria are triggered. The number of times Switching Criterion I is triggered is bounded by Lemma B.12. On the other hand, the number of times Switching Criterion II is triggered is equal to the number of times determinant of ${\mathbf{H}}_{t}$ doubles, which is bounded by Lemma B.17. Thus in total, the number of policy changes in Algorithm 2 is upper bounded by $2dR^{2}\kappa\gamma^{2}\log(ST/\delta)+\textbf{c}d\log(T)~{}.$
@@ B.6 Some Useful Lemmas
Lemma B.14 (Elliptic Potential Lemma (Lemma 10 [4])).
@@
Lemma B.15 (Lemma 12 of [4]).
@@
Lemma B.16 (Lemma 10 of [4]).
@@ Appendix C Useful Properties of GLMs
As a consequence, we have the following simple modification of the self-concordance results of [6].
@@
Putting $z=z_{2}$ establishes 30. To show 29, we further set $z=z_{1}+u(z_{2}-z_{1})$ for $u\in[0,1]$, (note that $z\geq z_{1}$) and integrate on $u$,
@@
For 31, we have, by application of 30, $\dot{\mu}(z_{1}+v(z_{2}-z_{1}))\geq\dot{\mu}(z_{1})\exp(R|v(z_{2}-z_{1})|)$. Therefore,
@@
$\geq\dot{\mu}(z_{1})\cdot\frac{1}{2+R|z_{1}-z_{2}|}$ (Lemma 10 of [2])
@@
In [8] Section 1.4.1, the author claims that if the GLM parameter $z$ lies in a bounded set, then the GLM is self-concordant, i.e., $|\ddot{\mu}(z)|\leq a\dot{\mu}(z)$, for some appropriate constant $a$ over this bounded set. Thereafter the author notes that the techniques developed in [8] guarantees $\kappa$-free regret rates (in $\sqrt{T}$ term) for such GLMs (i.e., all GLMs with bounded parameter). However, the claim regarding self-concordance of GLMs is not true in general. There are classes of GLMs whose parameters may be restricted in a bounded set, but for them no constant $a$ exists. One such example is the exponential distribution. The link function $\mu$ for exponential distribution is given as $\mu(z)=-\frac{1}{z}$. If we allow $z$ to lie in the set $(-c,0)$ for some positive $c$, then we have $\mu(z)$ strictly increasing (satisfying our assumption on monotonicity of $\mu$, thus a valid example). However, for this GLM,
@@ Appendix D Computational Cost
Consider a log-loss minimization oracle that returns the unconstrained MLE for a given GLM class with a computational complexity of $C_{opt}\cdot n$, when the log-loss is computed over $n$ data points. Let the maximum number of arms available every round be $K$. Furher, let the computational cost of an oracle that solves the non-convex optimization 9 be $NC_{opt}$.
@@
Comparison with ECOLog [7] shows that execution time for RS-GLinCB is significantly smaller. We posit that this is because RS-GLinCB solves a large convex optimization problem but less frequently, resulting into smaller overhead at the implementation level, while ECOLog solves a smaller convex optimization problem, but does so every round. On an implementation level, this translates into more function calls and computation. Further, we observe that with increasing $\kappa$, the execution time of RS-GLinCB increases, which is in accordance with Lemma 4.1 that quantifies the number of policy switches as an increasing function of $\kappa$.
While comparing with GLOC [14], we observe that RS-GLinCB performs better than GLOC in both high and low $\kappa$ regimes. Since GLOC runs an online convex optimization (online Newton step) algorithm to generate its confidence sets, the time taken by GLOC is nearly constant with changing $\kappa$. On the other hand, in accordance with Lemma 4.1, the computational cost of RS-GLinCB increases with $\kappa$. However, after a few initial rounds, when neither of the switching criteria are triggered, RS-GLinCB does not need to solve any computationally intensive optimization problem, hence these rounds execute very fast. In practice, with typical data distribution, RS-GLinCB reaches this stage much before what the worst-case guarantees show, hence we see it perform better than GLOC.
@@ Appendix E Projection
We describe the projection step used in Algorithms 1 and 2. We present arguments similar to the ones made in Appendix B.3 of [6]. We write
@@
$\widetilde{\theta}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}\left\lVert\sum_{s=1}^{t}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu\left(\langle x_{s},\widehat{\theta}\rangle\right)\right)x_{s}-\lambda\widehat{\theta}\right\rVert_{{\mathbf{H}}(\theta)^{-1}}$ (32)
@@
$\left\lVert\widetilde{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}(\theta^{*})}\leq 2(1+c)\left\lVert\sum_{s=1}^{t}(\mu\left(\langle x_{s},\theta^{*}\rangle\right)-r_{s})x_{s}\right\rVert_{{\mathbf{H}}(\theta^{*})^{-1}}+2S\sqrt{\lambda(1+c)}$
@@
Let us define the matrix ${\mathbf{G}}=\sum_{s\in[t]}\alpha(x,\widetilde{\theta},\theta^{*})x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$. Using the above fact, we obtain the relation: ${\mathbf{G}}\succeq\frac{1}{1+c}{\mathbf{H}}^{*}$ and ${\mathbf{G}}\succeq\frac{1}{1+c}{{\mathbf{H}}}(\widetilde{\theta})$. Now,
@@
$=\sqrt{1+c}\left\lVert\sum_{s\in[t]}\left(\alpha(x_{s},\widetilde{\theta},\theta^{*})\langle\widetilde{\theta}-\theta^{*},x_{s}\rangle\right)x_{s}+\lambda(\widetilde{\theta}-\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}$
$=\sqrt{1+c}\left\lVert\sum_{s\in[t]}\left(\mu\left(\langle x_{s},\widetilde{\theta}\rangle\right)-\mu\left(\langle x_{s},\theta^{*}\rangle\right)\right)x_{s}+\lambda(\widetilde{\theta}-\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}$ (Taylor’s theorem)
$=\sqrt{1+c}\left\lVert\left(\sum_{s\in[t]}\mu\left(\langle\widetilde{\theta},x_{s}\rangle\right)x_{s}+\lambda\widetilde{\theta}\right)-\left(\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}+\lambda\theta^{*}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
Let $g(\theta)=\sum_{s=1}^{t}\mu\left(\langle x_{s},\theta\rangle\right)x_{s}+\lambda\theta$ for any $\theta$. Therefore, we have,
@@
$\leq\sqrt{1+c}\left(\left\lVert g(\widetilde{\theta})-g(\widehat{\theta})+g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}\right)$
@@
$\leq\sqrt{1+c}\left(\left\lVert\sum_{s\in[t]}\mu\left(\langle\widetilde{\theta},x_{s}\rangle\right)x_{s}-g(\widehat{\theta})\right\rVert_{{\mathbf{G}}^{-1}}+\left\lVert g(\widehat{\theta})-\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}\right\rVert_{{\mathbf{G}}^{-1}}\right)+2S\sqrt{\lambda(1+c)}$ ($\triangle$ inequality)
$\leq(1+c)\left(\left\lVert\sum_{s\in[t]}\mu\left(\langle\widetilde{\theta},x_{s}\rangle\right)x_{s}-g(\widehat{\theta})\right\rVert_{{{\mathbf{H}}(\widetilde{\theta})}^{-1}}+\left\lVert g(\widehat{\theta})-\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}\right)+2S\sqrt{\lambda(1+c)}$ (${\mathbf{H}}^{*-1}\succeq({1+c}){\mathbf{G}}^{-1}$)
$\leq 2(1+c)\left\lVert g(\widehat{\theta})-\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}+2S\sqrt{\lambda(1+c)}$ (by (32))
$=2(1+c)\left\lVert\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}-\sum_{s\in[t]}r_{s}x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}+2S\sqrt{\lambda(1+c)}$ ($\widehat{\theta}$ is the unconstrained MLE, $g(\widehat{\theta})=\sum_{s\in[t]}r_{s}x_{s}+\lambda\widehat{\theta}$.)